If an economy consumes 75% of any increase in income, then an increase in autonomous investment of $1 billion could result in an increase in RGDP of as much as:

a. $4.0 billion

b. $5.0 billion

c. $1.8 billion

d. $6.0 billion

e. $1.0 billion

Please show all work and as simple as possible!
Answer #1

Answer

option a

-----

A 75% consumption of disposable income means a person consumes $0.75 from $1 income.

It means the marginal propensity to consume(MPC) is 0.75.

Formula for the multiplier is:

Multiplier =1/(1-MPC)

=1/(1-0.75)

=4

the maximum change in GDP=change in autonomous investment * multiplier

=1*4

=$4 billion

increase RGDP of as much as $4 billion